# CITYWALK: Enhancing LLM-Based C++ Unit Test Generation via Project-Dependency Awareness and Language-Specific Knowledge

### v1

@@ 2.6. Unit Test Generation
Unit test generation is a complex task that poses significant challenges when generating high-quality test cases from scratch with limited guidance. To address this, CITYWALK initially produces a test suite for the given focal method by following step-by-step instructions. Subsequently, CITYWALK corrects the failed test cases using a three-phase rule-based fixing approach. As illustrated in Algorithm 1, we provide detailed overview of the unit test generation process as follows.
@@ 2.6.1. Initial Unit Test Case Generation via Step-by-Step Instructions
    Unit Test Generation (Line 5): Based on the $\mathtt{Intent}$ of $\mathtt{m_{focal}}$, the LLM generates an initial test suite for $\mathtt{m_{focal}}$ by utilizing the provided contextual information, which includes the focal contexts $\mathtt{Context_{focal}}$, project dependencies $\mathtt{Dep_{c}}$ and $\mathtt{Dep_{d}}$, intention contexts $\mathtt{Context_{intent}}$. Additionally, the LLM matches the contextual information with the extracted $\mathtt{Dep_{ingredient}}$ to ensure an precise understanding of the purpose of $\mathtt{m_{focal}}$.
@@
    Test Case Refinement (Line 7): To enhance the effectiveness of the LLM-generated test cases $\mathtt{TC}$, the LLM refines the initial test suite using the language-specific domain knowledge as guidelines $\mathtt{Guideline_{DK}}$, ensuring the generation of high coverage and executable $\mathtt{TC}$.
@@ 2.6.2. Post-Processing of the LLM-Generated Unit Test Cases
    • If the project uses gtest, retaining only one main function in the generated test suite; otherwise, ensuring each test method is called by the main function.
@@
  3. LLM-based Compilation Error Fixing: If compilation errors persist after the previous phase, CITYWALK transitions to a one-round LLM-based fixing phase (Lines 12-14). During this phase, CITYWALK collects details about the failed test cases along with their associated compilation error messages to construct the fixing prompt. The LLM is then prompted to analyze the root cause of the errors and make corrections to the failed test cases. If any test case continues to fail compilation after this phase, it will be removed from $\mathtt{TC}$.
@@ 3.1. Research Questions
To assess the effectiveness of CITYWALK, we raise the following three research questions (RQs):
  • RQ1: How does CITYWALK perform in C++ unit test generation when compared to state-of-the-art baselines? This RQ aims to evaluate the superior effectiveness of CITYWALK in comparison to open-source code LLMs, closed-source general-purpose LLMs, and LLM-based unit test generation approaches within the context of C++ unit test generation. To achieve this, we conduct a comprehensive evaluation of CITYWALK against six baselines using a diverse collection of C++ projects.
@@
  • RQ3: What is the generalizability of CITYWALK to additional LLMs? This RQ assesses the generalization capabilities of CITYWALK by extending it to two open-source LLMs tailored to the code generation task, thereby enhancing evaluation diversity.
@@
Project Application Type GitHub Stars Size (LoC) # of Files # of Focal Methods # of Complex Methods Trained?
hjson-cpp  [^4] https://github.com/hjson/hjson-cpp User Interface for JSON 70 2911 4 25 9 (36.0%) Y
tinyxml2  [^5] https://github.com/leethomason/tinyxml2 XML Parser 5.2K 3606 1 158 5 (3.2%) Y
yaml-cpp  [^6] https://github.com/jbeder/yaml-cpp YAML Parser and Emitter 5.3K 8800 28 204 31 (15.2%) Y
re2  [^7] https://github.com/google/re2 Regular Expression Engine 9.1K 20373 10 146 41 (28.1%) Y
ninja  [^8] https://github.com/ninja-build/ninja Build System 11.5K 37512 19 238 38 (16.0%) Y
leveldb  [^9] https://github.com/google/leveldb Key-Value Storage Library 37K 149371 17 220 6 (2.7%) Y
json.cpp  [^10] https://github.com/jart/json.cpp JSON Parsing Library 689 62677 1 34 4 (11.8%) N
glomap  [^11] https://github.com/colmap/glomap Map Management Library 1.6K 8477 9 32 4 (12.5%) N
3.2. Benchmark
To comprehensively evaluate the quality of LLM-generated test cases, we construct a new benchmark consisting of eight real-world open-source C++ projects from GitHub. As shown in Table 2, the first four projects are widely utilized in recent studies on automated C++ unit testing (Herlim et al., 2022; Rho et al., 2023, 2024). Additionally, we crawl two practical projects related to basic software: ninja (build system) and leveldb (database). To address potential data leakage concerns, we include two projects—json.cpp and glomap—which were created after the knowledge cutoff date (October 2023). This ensures that existing LLMs have not been trained on these projects. The selection criteria for these projects are as follows. First, each project has received more than 50 stars on GitHub and is actively maintained, ensuring ongoing community interest and updates. Second, the projects span a range of application domains, including parsers (e.g., XML or YAML), basic software systems, and algorithm libraries. Third, the projects feature a number of complex methods with a cyclomatic complexity (McCabe, 1976) greater than 10, which suggests the presence of nested control flows, while also covering intricate C++ language features. The size of the selected projects varies from 2.9K lines of code (LoC) to 149.4K LoC. Larger projects are excluded to effectively manage token costs within our limited budget. In total, CITYWALK generates C++ unit test cases for 1057 focal methods across these eight projects. These selection criteria ensure that our benchmark is both high-quality and reproducible, while maintaining a balance between diversity and resource constraints.
@@ 3.3. Baselines
This paper focuses on addressing the C++ unit test generation task using LLMs. To this end, we compare CITYWALK against six state-of-the-art baselines. First, we select two representative open-source code LLMs: CodeGeeX4 (Zheng et al., 2023) and DeepSeek-V2.5 (DeepSeek-AI, 2024b, a), both of which have demonstrated competitive performance on recent public benchmarks related to coding tasks, such as BigCodeBench (Zhuo et al., 2024) and NaturalCodeBench (Zhang et al., 2024b). The selection of these two models is motivated by their distinct characteristics in terms of parameter size (CodeGeeX4 has 9B parameters, while DeepSeek-V2.5 comprises 236B parameters) and architecture (CodeGeeX4 utilizes a Transformer-based architecture, whereas DeepSeek-V2.5 employs a Mixture of Experts-based architecture). We exclude CodeLLaMA (Rozière et al., 2023) from our evaluation due to its limited context length. Additionally, we include the closed-source general-purpose LLMs (i.e., ChatGPT (OpenAI, 2022) and GPT-4o (OpenAI, 2023)) because of their established effectiveness across a broad spectrum of tasks. Finally, we compare CITYWALK with two LLM-based unit test generation approaches for Java: ChatTester (Yuan et al., 2024) and HITS (Wang et al., 2024b). Since existing automated C++ unit test generation tools, such as Coyote (Rho et al., 2023) and CITRUS (Herlim et al., 2022), are currently unavailable, we are unable to reproduce the results reported in their respective papers. Therefore, we do not include them as baselines in this paper.
@@ 4.1. Answering RQ1
To answer this question, we conduct a comprehensive comparison of CITYWALK against six state-of-the-art baselines using the collected benchmark. For each LLM baseline, we utilize its inference API for implementation, employing the same basic prompt described in Section 1. For ChatTester (Yuan et al., 2024) and HITS (Wang et al., 2024b), we use their open-source reproduction artifacts for implementation. Furthermore, we apply the same configuration settings as those used in CITYWALK for fair comparison.
@@
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg.
CodeGeeX4 0.00% 6.95% 38.98% 0.09% 4.22% 11.75% 10.70% 35.75% 13.56%
DeepSeek-V2.5 0.00% 0.00% 17.79% 0.00% 7.35% 13.55% 25.14% 51.93% 14.47%
ChatGPT 0.00% 0.00% 19.34% 5.45% 5.00% 9.64% 31.16% 28.57% 12.40%
GPT-4o 16.17% 41.26% 22.01% 0.00% 6.90% 9.74% 51.62% 49.70% 24.68%
ChatTester 20.00% 18.02% 6.42% 4.68% 11.36% 0.00% 18.75% 8.82% 11.01%
HITS 1.10% 5.94% 5.65% 0.00% 20.37% 2.66% 0.00% 1.92% 4.71%
CITYWALK 91.84% 97.25% 65.01% 64.85% 47.34% 47.49% 100.00% 47.65% 70.18%
@@ 4.1.1. Experimental Metric Evaluation
    CITYWALK demonstrates superior performance compared to state-of-the-art baselines on the collected benchmark. When compared to the six LLM-based baselines, CITYWALK achieves the best results across all evaluation metrics (except CSR and EPR in the glomap project). A closer examination of the results reveals that the second-best outcomes for each metric vary depending on the project under test. Nevertheless, GPT-4o consistently outperforms other baselines regarding average scores of all the four evaluation metrics (listed in the Avg. rows). Specifically, CITYWALK surpasses the best baseline GPT-4o by 42.05%, 32.86%, 26.41%, and 23.09% in terms of CSR, ERP, $\mathbf{Cov_{L}}$, and $\mathbf{Cov_{B}}$, respectively. These improvements underscore the effectiveness of CITYWALK in the C++ unit test generation task.
    Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg.
    CodeGeeX4 0.00% 6.00% 2.76% 0.09% 3.82% 0.36% 8.96% 15.54% 4.69%
    DeepSeek-V2.5 0.00% 0.00% 13.78% 0.00% 6.69% 12.02% 24.02% 39.91% 12.05%
    ChatGPT 0.00% 0.00% 18.37% 3.41% 4.04% 8.66% 30.82% 10.48% 9.47%
    GPT-4o 14.37% 38.46% 15.75% 0.00% 6.57% 7.61% 49.10% 47.88% 22.47%
    ChatTester 12.00% 13.95% 4.91% 1.75% 8.77% 0.00% 9.38% 8.82% 7.45%
    HITS 1.10% 5.45% 4.52% 0.00% 14.90% 1.73% 0.00% 1.92% 3.70%
    CITYWALK 57.14% 88.24% 51.03% 49.37% 42.47% 41.04% 98.91% 38.26% 58.31%
    Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg.
    CodeGeeX4 0.00% 15.30% 16.54% 0.00% 8.44% 10.45% 7.14% 21.75% 9.95%
    DeepSeek-V2.5 0.00% 0.00% 31.50% 0.00% 20.94% 22.23% 37.31% 30.04% 17.75%
    ChatGPT 0.00% 0.00% 25.59% 0.84% 10.68% 14.73% 12.09% 16.75% 10.09%
    GPT-4o 7.73% 43.21% 30.17% 0.00% 10.64% 15.45% 26.50% 22.31% 19.50%
    ChatTester 1.68% 45.00% 0.00% 3.05% 17.52% 0.00% 3.56% 4.47% 9.41%
    HITS 0.00% 35.64% 25.56% 0.00% 19.55% 8.53% 0.00% 4.15% 11.68%
    CITYWALK 27.94% 77.61% 44.46% 42.56% 56.60% 25.25% 50.71% 42.15% 45.91%
@@
    Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg.
    CodeGeeX4 0.00% 9.97% 15.73% 0.00% 9.18% 7.88% 10.81% 22.22% 9.47%
    DeepSeek-V2.5 0.00% 0.00% 24.24% 0.00% 19.03% 23.24% 32.58% 28.89% 16.00%
    ChatGPT 0.00% 0.00% 21.61% 0.51% 10.22% 12.50% 13.06% 13.33% 8.90%
    GPT-4o 8.82% 37.18% 23.54% 0.00% 8.97% 14.06% 24.84% 17.78% 16.90%
    ChatTester 0.89% 38.45% 0.00% 2.24% 15.46% 0.00% 1.61% 2.48% 7.64%
    HITS 0.00% 26.74% 17.70% 0.00% 16.46% 8.72% 0.00% 3.09% 9.09%
    CITYWALK 24.29% 65.66% 35.56% 35.64% 49.72% 24.64% 46.61% 37.82% 39.99%
@@
    CITYWALK effectively generates fewer compilation errors than the selected baselines. As observed from Table 3, it is important to that, in many instances, the selected baselines fail to generate any compilable unit test cases for the projects under test using the basic prompt. This issue arises because projects such as hjson-cpp, tinyxml2, re2, and json.cpp do not include the gtest testing framework. Consequently, when the LLMs generate test code without accounting for these configuration dependencies, they default to using the gtest framework, resulting in compilation errors for all generated test cases. In contrast, the higher CSR achieved by CITYWALK is largely due to the inclusion of additional project dependencies. CITYWALK guides the LLMs to generate syntactically correct test code by directly incorporating relevant configuration dependency information into the prompt, thereby greatly reducing the occurrence of compilation errors.
    Table 7. Comparison of CITYWALK against the Baselines in Terms of the Number and Complexity of Generated Test Cases
    Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg. Num Avg. Size Avg. CC
    CodeGeeX4 175 417 652 1133 1231 1115 402 193 664.8 7.4 1.1
    DeepSeek-V2.5 463 1405 1147 1466 2273 1572 533 233 1136.5 7.5 1.1
    ChatGPT 59 172 827 587 1139 1235 292 105 552.0 6.8 1.1
    GPT-4o 167 143 750 777 1218 986 277 165 560.4 8.1 1.2
    ChatTester 25 172 265 171 308 163 32 34 146.3 7.1 1.3
    HITS 181 404 1239 1190 1389 752 233 156 693.0 6.0 1.0
    CITYWALK 49 255 243 219 657 386 183 123 264.4 9.7 1.5
@@
    CITYWALK achieves high code coverage with fewer test cases than the baselines. Table 7 reports the number of test cases generated by CITYWALK and the baselines for each project. The Avg. Num represents the average number of generated test cases across all evaluated projects, while the Avg. Size and Avg. CC represent the average source code lines and cyclomatic complexity of all generated test cases, respectively. The statistical results indicate that CITYWALK generates significantly fewer test cases compared to the selected baselines (except ChatTester), proving that CITYWALK does not rely on sampling a large number of test cases to achieve advantages. The reason ChatTester generates the fewest test cases lies in its prompt design, which explicitly instructs the LLM to “write one test case for each focal method”. This constraint significantly limits the number of test cases generated by ChatTester, resulting in suboptimal performance of ChatTester across many projects. Meanwhile, the scale and complexity of the test cases generated by CITYWALK are slightly higher than those of the LLMs. This outcome is reasonable, as generating correct test assertions necessitates a certain level of complex code logic.
@@
    The correctness and coverage metrics do not necessarily indicate a positive correlation. As illustrated in Tables 3, Tables 4, Tables 5 and Tables 6, high correctness does not always guarantee high code coverage, and low correctness can sometimes lead to higher coverage. For instance, DeepSeek-V2.5 on yaml-cpp demonstrates this phenomenon. Despite achieving lower CSR and EPR compared to other LLM baselines, DeepSeek-V2.5 attains the second-highest code coverage scores. According to the fourth column in Table 7, it is evident that DeepSeek-V2.5 generates more test cases for yaml-cpp than other LLM baselines. Consequently, the proportion of incorrect test cases is also relatively high, leading to a scenario characterized by low correctness but high coverage.

### v2

@@ 2.6. Unit Test Generation
Unit test generation is a complex task that poses significant challenges when generating high-quality test cases from scratch with limited guidance. To address this, CITYWALK initially produces an entire test file for the given focal method by following step-by-step instructions. Subsequently, CITYWALK corrects the failed test cases using a three-phase rule-based fixing approach. As illustrated in Algorithm 1, we provide detailed process overview as follows.
@@ 2.6.1. Initial Unit Test Case Generation via Step-by-Step Instructions
    Unit Test Generation (Line 5): Based on the $\mathtt{Intent}$ of $\mathtt{m_{focal}}$, the LLM generates an initial test file for $\mathtt{m_{focal}}$ by utilizing the provided contextual information, which includes the focal contexts $\mathtt{Context_{focal}}$, project dependencies $\mathtt{Dep_{c}}$ and $\mathtt{Dep_{d}}$, intention contexts $\mathtt{Context_{intent}}$. Additionally, the LLM matches the contextual information with the extracted $\mathtt{Dep_{ingredient}}$ to ensure an precise understanding of the purpose of $\mathtt{m_{focal}}$.
@@
    Test Case Refinement (Line 7): To enhance the effectiveness of the LLM-generated test cases $\mathtt{TC}$, the LLM refines the initial test file using the language-specific domain knowledge as guidelines $\mathtt{Guideline_{DK}}$, ensuring the generation of high coverage and executable $\mathtt{TC}$.
@@ 2.6.2. Post-Processing of the LLM-Generated Unit Test Cases
    • If the project uses gtest, retaining only one main function in the generated test file; otherwise, ensuring each test method is called by the main function.
@@
  3. LLM-based Compilation Error Fixing: If compilation errors persist after Phase ➋, CITYWALK transitions to a one-round LLM-based fixing phase (Lines 12-14). During this phase, CITYWALK collects details about the failed test cases along with their associated compilation error messages to construct the fixing prompt. The LLM is then prompted to analyze the root cause of the errors and make corrections to the failed test cases. If any test case continues to fail compilation after this phase, it will be removed from $\mathtt{TC}$.
@@ 3.1. Research Questions
To assess the effectiveness of CITYWALK, we raise the following two research questions (RQs):
  • RQ1: How does CITYWALK perform in C++ unit test generation when compared to state-of-the-art baselines? This RQ aims to evaluate the superior effectiveness of CITYWALK in comparison to open-source code LLMs, closed-source general-purpose LLMs, and LLM-based unit test generation approaches within the context of C++ unit test generation. To achieve this, we conduct a comprehensive evaluation of CITYWALK against seven baselines using a collection of ten C++ projects. Furthermore, we assess CITYWALK’s generalization capabilities using three additional LLM baselines, thereby enhancing evaluation diversity.
@@
3.2. Benchmark
To comprehensively evaluate the quality of LLM-generated test cases, we construct a new benchmark consisting of ten real-world open-source C++ projects from GitHub. As shown in Table 3, the first four projects are widely utilized in recent studies on automated C++ unit testing (Herlim et al., 2022; Rho et al., 2023, 2024). Additionally, we crawl two practical projects related to basic software: ninja (build system) and leveldb (database). To mitigate potential data leakage concerns, we include four projects—json.cpp, glomap, papy, and mlx—all created after the GPT-4o knowledge cutoff date (October 2023). This ensures that GPT-4o was not trained on these projects. The selection criteria for these projects are as follows. First, each project has received more than 50 stars on GitHub and is actively maintained, ensuring ongoing community interest and updates. Second, the projects span a range of application domains, including parsers (e.g., XML or YAML), basic software systems, and algorithm libraries. Third, the projects feature a number of complex methods with a cyclomatic complexity (McCabe, 1976) greater than 10, which suggests the presence of nested control flows, while also covering intricate C++ language features. The size of the selected projects varies from 1.9K lines of code (LoC) to 149.4K LoC. Larger projects are excluded to effectively manage token costs within our limited budget. In total, CITYWALK generates C++ unit test cases for 1288 focal methods across these ten projects. These selection criteria ensure that our benchmark is both high-quality and reproducible, while maintaining a balance between diversity and resource constraints.
@@ 3.2. Benchmark
Project Application Type GitHub Stars Size (LoC) # Files # Focal Methods # Complex Methods Trained?
hjson-cpp [^4] https://github.com/hjson/hjson-cpp User Interface for JSON 73 2911 4 25 9 (36.0%) Y
tinyxml2555https://github.com/leethomason/tinyxml2 XML Parser 5.4K 3606 1 158 5 (3.2%) Y
yaml-cpp [^6] https://github.com/jbeder/yaml-cpp YAML Parser and Emitter 5.6K 8800 28 204 31 (15.2%) Y
re2777https://github.com/google/re2 Regular Expression Engine 9.4K 20373 10 146 41 (28.1%) Y
ninja [^8] https://github.com/ninja-build/ninja Build System 12.2K 37512 19 238 38 (16.0%) Y
leveldb [^9] https://github.com/google/leveldb Key-Value Storage Library 38K 149371 17 220 6 (2.7%) Y
json.cpp [^10] https://github.com/jart/json.cpp JSON Parsing Library 747 62677 1 34 4 (11.8%) N
glomap [^11] https://github.com/colmap/glomap Map Management Library 1.9K 8477 9 32 4 (12.5%) N
papy [^12] https://github.com/noahpop77/Papy JSON Data Generator 62 1869 6 25 2 (8.0%) N
mlx [^13] https://github.com/ml-explore/mlx Array Framework 21.9K 20137 11 206 18 (8.7%) N
@@ 3.3. Baselines
This paper focuses on addressing the C++ unit test generation task using LLMs. To this end, we compare CITYWALK against seven state-of-the-art baselines. First, we select two representative open-source code LLMs: CodeGeeX4 (Zheng et al., 2023) and DeepSeek-V3 (DeepSeek-AI, 2024), both of which have demonstrated competitive performance on recent public benchmarks related to coding tasks, such as BigCodeBench (Zhuo et al., 2024) and NaturalCodeBench (Zhang et al., 2024b). The selection of these two models is motivated by their distinct characteristics in terms of parameter size (CodeGeeX4 has 9B parameters, while DeepSeek-V3 comprises 671B parameters) and architecture (CodeGeeX4 utilizes a Transformer-based architecture, whereas DeepSeek-V3 employs a Mixture of Experts-based architecture). We exclude CodeLLaMA (Rozière et al., 2023) from our evaluation due to its limited context length. Additionally, we include the closed-source general-purpose LLMs (i.e., GPT-3.5 (OpenAI, 2022) and GPT-4o (OpenAI, 2023)) because of their established effectiveness across a broad spectrum of tasks. Finally, we compare CITYWALK with three LLM-based unit test generation approaches for Java and JavaScript: ChatTester (Yuan et al., 2024), HITS (Wang et al., 2024b), and TestPilot (Schäfer et al., 2024). Since existing automated C++ unit test generation tools, such as Coyote (Rho et al., 2023) and CITRUS (Herlim et al., 2022), are currently unavailable, we are unable to reproduce the results reported in their respective papers. Therefore, we do not include them as baselines in this paper.
@@ 4.1. Answering RQ1
To answer this question, we conduct a comprehensive comparison of CITYWALK against seven baselines using the collected benchmark. For each LLM, we utilize its inference API for implementation, employing the same basic prompt described in Section 1. Specifically, we employ the gpt-3.5-turbo-0125 version of the GPT-3.5 model. For ChatTester (Yuan et al., 2024), HITS (Wang et al., 2024b), and TestPilot (Schäfer et al., 2024), our implementation is based on their open-source reproduction artifacts from GitHub. Furthermore, we apply the same configuration settings as those used in CITYWALK for fair comparison [^14] Note that the maximum prompt length used in our experiments (14733 tokens) remains within the context window limits of all evaluated LLMs—GPT-3.5 (16K), CodeGeeX4 (128K), DeepSeek-V3 (128K), and GPT-4o (128K). Therefore, no prompt truncation is necessary during any of the experiments..
@@
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap papy mlx Avg.
CodeGeeX4 0.00% 6.95% 38.98% 0.09% 4.22% 11.75% 10.70% 35.75% 50.00% 51.35% 20.98%
DeepSeek-V3 0.00% 0.00% 17.79% 0.00% 7.35% 13.55% 25.14% 51.93% 18.75% 68.34% 20.29%
GPT-3.5 0.00% 0.00% 19.34% 5.45% 5.00% 9.64% 31.16% 28.57% 52.22% 72.80% 22.42%
GPT-4o 16.17% 41.26% 22.01% 0.00% 6.90% 9.74% 51.62% 49.70% 47.90% 73.07% 31.84%
ChatTester 20.00% 18.02% 6.42% 4.68% 11.36% 0.00% 18.75% 8.82% 20.00% 18.01% 12.61%
HITS 1.10% 5.94% 5.65% 0.00% 20.37% 2.66% 0.00% 1.92% 0.00% 0.00% 3.76%
TestPilot 26.67% 69.23% 20.47% 0.00% 14.22% 26.18% 70.68% 28.00% 54.17% 2.66% 31.23%
CITYWALK 100.00% 97.25% 80.28% 64.85% 70.09% 47.83% 100.00% 96.70% 84.29% 92.61% 83.39%
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap papy mlx Avg.
CodeGeeX4 0.00% 6.00% 2.76% 0.09% 3.82% 0.36% 8.96% 15.54% 40.54% 41.89% 12.00%
DeepSeek-V3 0.00% 0.00% 13.78% 0.00% 6.69% 12.02% 24.02% 39.91% 18.75% 66.29% 18.15%
GPT-3.5 0.00% 0.00% 18.37% 3.41% 4.04% 8.66% 30.82% 10.48% 38.89% 69.50% 18.42%
GPT-4o 14.37% 38.46% 15.75% 0.00% 6.57% 7.61% 49.10% 47.88% 46.22% 69.94% 29.59%
ChatTester 12.00% 13.95% 4.91% 1.75% 8.77% 0.00% 9.38% 8.82% 17.14% 13.27% 9.00%
HITS 1.10% 5.45% 4.52% 0.00% 14.90% 1.73% 0.00% 1.92% 0.00% 0.00% 2.96%
TestPilot 6.67% 64.55% 18.14% 0.00% 13.33% 23.98% 70.30% 12.00% 34.72% 2.66% 24.64%
CITYWALK 77.50% 88.24% 67.82% 49.37% 62.12% 41.00% 98.91% 78.02% 81.43% 89.13% 73.35%
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap papy mlx Avg.
CodeGeeX4 0.00% 15.30% 16.54% 0.00% 8.44% 10.45% 7.14% 21.75% 7.22% 8.11% 9.50%
DeepSeek-V3 0.00% 0.00% 31.50% 0.00% 20.94% 22.23% 37.31% 30.04% 27.78% 16.80% 18.66%
GPT-3.5 0.00% 0.00% 25.59% 0.84% 10.68% 14.73% 12.09% 16.75% 22.10% 17.60% 12.04%
GPT-4o 7.73% 43.21% 30.17% 0.00% 10.64% 15.45% 26.50% 22.31% 22.89% 17.85% 19.68%
ChatTester 1.68% 45.00% 0.00% 3.05% 17.52% 0.00% 3.56% 4.47% 9.93% 4.58% 8.98%
HITS 0.00% 35.64% 25.56% 0.00% 19.55% 8.53% 0.00% 4.15% 7.02% 0.00% 9.34%
TestPilot 0.00% 70.90% 0.00% 0.00% 0.00% 24.33% 48.17% 0.00% 22.11% 4.55% 17.01%
CITYWALK 27.94% 77.61% 44.46% 42.56% 56.60% 25.25% 50.71% 42.15% 53.11% 24.20% 44.46%
@@ 4.1.1. Experimental Metric Evaluation
    CITYWALK demonstrates superior performance compared to state-of-the-art baselines on the collected benchmark. When compared to the seven baselines, CITYWALK achieves the best results across all evaluation metrics. A closer examination of the results reveals that the second-best outcomes for each metric vary depending on the project under test. Nevertheless, GPT-4o consistently outperforms other baselines regarding average scores (listed in the Avg. rows) of all the four evaluation metrics. Specifically, CITYWALK surpasses the best baseline, GPT-4o, by 51.55% in CSR, 43.76% in ERP, 24.78% in $\mathbf{Cov_{L}}$, and 21.55% in $\mathbf{Cov_{B}}$. These improvements underscore the effectiveness of CITYWALK in the C++ unit test generation task.
@@
    Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap papy mlx Avg.
    CodeGeeX4 0.00% 9.97% 15.73% 0.00% 9.18% 7.88% 10.81% 22.22% 1.35% 2.70% 7.98%
    DeepSeek-V3 0.00% 0.00% 24.24% 0.00% 19.03% 23.24% 32.58% 28.89% 15.78% 13.20% 15.70%
    GPT-3.5 0.00% 0.00% 21.61% 0.51% 10.22% 12.50% 13.06% 13.33% 8.82% 13.32% 9.34%
    GPT-4o 8.82% 37.18% 23.54% 0.00% 8.97% 14.06% 24.84% 17.78% 12.87% 13.40% 16.15%
    ChatTester 0.89% 38.45% 0.00% 2.24% 15.46% 0.00% 1.61% 2.48% 4.23% 2.27% 6.76%
    HITS 0.00% 26.74% 17.70% 0.00% 16.46% 8.72% 0.00% 3.09% 0.00% 0.00% 7.27%
    TestPilot 0.00% 57.14% 0.00% 0.00% 0.00% 21.48% 45.48% 0.00% 8.83% 2.27% 13.52%
    CITYWALK 24.29% 65.66% 35.56% 35.64% 49.72% 24.64% 46.61% 37.82% 38.97% 18.05% 37.70%
    Table 8. Comparison of CITYWALK against the Baselines in Terms of the Number of Generated Test Cases
    Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap papy mlx Total Failed
    CodeGeeX4 175 417 652 1133 1231 1115 402 193 74 439 5831 5456 (93.57%)
    DeepSeek-V3 463 1405 1147 1466 2273 1572 533 233 160 482 9734 8664 (89.01%)
    GPT-3.5 59 172 827 587 1139 1235 292 105 90 458 4964 4185 (84.30%)
    GPT-4o 167 143 750 777 1218 986 277 165 119 479 5081 4124 (81.16%)
    ChatTester 25 172 265 171 308 163 32 34 35 211 1416 1306 (92.23%)
    HITS 181 404 1239 1190 1389 752 233 156 78 1033 6655 6352 (95.45%)
    TestPilot 15 299 215 94 225 955 266 25 72 188 2354 1642 (69.75%)
    CITYWALK 40 255 289 140 565 922 183 91 70 230 2785 1021 (36.66%)
@@
    CITYWALK effectively generates fewer compilation errors than the selected baselines. As observed from Table 4, it is important to that, in many instances, the selected baselines fail to generate any compilable unit test cases for the projects under test using the basic prompt. This issue arises because projects (e.g., hjson-cpp) do not include the gtest testing framework. Consequently, when the LLMs generate test code without accounting for these configuration dependencies, they default to using the gtest framework, resulting in compilation errors for all generated test cases. If configuration-dependent information is incorporated into other baseline approaches, a portion of their compilation errors could be resolved and their performance could potentially be improved. In contrast, the higher CSR achieved by CITYWALK is largely due to the inclusion of additional project dependencies. CITYWALK guides the LLMs to generate syntactically correct test code by directly incorporating relevant configuration dependency information into the prompt, thereby greatly reducing the occurrence of compilation errors.
@@
    CITYWALK achieves high code coverage with fewer test cases than the baselines. Table 8 reports the number of test cases generated by CITYWALK and the baselines for each project. The Total column represents the overall number of test cases generated across all projects. The statistical results indicate that CITYWALK generates fewer test cases compared to the selected baselines (except ChatTester and TestPilot), proving that CITYWALK does not rely on sampling a large number of test cases to achieve advantages. The reason ChatTester generates the fewest test cases lies in its prompt design, which explicitly instructs the LLM to “write one test case for each focal method”. This constraint significantly limits the number of test cases generated by ChatTester, resulting in suboptimal performance of ChatTester across many projects. Furthermore, we examine the number of failed test cases generated by CITYWALK and the baselines, as reported in the Failed column of Table 8. The results show that CITYWALK produces significantly fewer failed test cases compared to the baselines, with only 36.66% of its generated test cases failing to compile or execute. This finding suggests that CITYWALK is more effective in generating syntactically and semantically correct test cases while maintaining high code coverage.
@@
    Comparison of the code complexity and mock usage between CITYWALK and the baselines. As shown in Table 9, the rows Avg. Size and Avg. CC represent the average number of source code lines and the average cyclomatic complexity of the generated test cases, respectively. Overall, the test cases generated by CITYWALK exhibit slightly greater scale and complexity compared to the baselines. This is expected, as generating correct test assertions necessitates a certain level of complex code logic. In addition, we examine the frequency of mock usage, defined as the proportion of test files that employ mock objects. CITYWALK demonstrates a significantly higher mock usage frequency of 41.73%, indicating its greater capability in producing test cases that involve mocking. This is particularly important for effectively testing complex methods and achieving comprehensive code coverage. Beyond the quantitative analysis, we assess the adherence of the generated mocks to established best practices (Spadini et al., 2017). Figure 7 presents a CITYWALK-generated test case for Plan::AddSubTarget from the ninja project. In this example, the core logic of the Plan module is retained, while its dependency nodes and edges are replaced with mocks to simulate various conditions such as dirty or ready states. This strategy enables precise and isolated testing without compromising behavioral integrity or introducing excessive mocking.
    Table 9. Comparison of CITYWALK against the Baselines in Terms of the Complexity and Mock Frequency of Generated Test Cases
    Approach CodeGeeX4 DeepSeek-V3 GPT-3.5 GPT-4o ChatTester HITS TestPilot CITYWALK
    Avg. Size 7.7 7.4 6.8 8.0 7.3 5.9 7.1 9.5
    Avg. CC 1.3 1.1 1.1 1.2 1.3 1.0 1.4 1.5
    Mock Frequency 26.93% 25.38% 5.60% 25.15% 15.23% 21.44% 4.19% 41.73%
    Figure 7. Illustrative Test Case with Compliant Mock Usage.
[…]
  6. (6)
    The correctness and coverage metrics do not necessarily indicate a positive correlation. As illustrated in Table 4, Table 5, Table 6, and Table 7, high correctness does not always guarantee high code coverage, and low correctness can sometimes lead to higher coverage. For instance, DeepSeek-V3 on yaml-cpp demonstrates this phenomenon. Despite achieving lower CSR and EPR compared to other LLM baselines, DeepSeek-V3 attains the second-highest code coverage scores. According to the fourth column in Table 8, it is evident that DeepSeek-V3 generates more test cases for yaml-cpp than other LLM baselines. Consequently, the proportion of incorrect test cases is also relatively high, leading to a scenario characterized by low correctness but high coverage.